OpenStudy (anonymous):

i qot a new one ppl . findd the vaalue ovv x for whichh x dividded bii 7 andd 7 dividded bii x are equaal

OpenStudy (anonymous):

>.<

OpenStudy (anonymous):

7

OpenStudy (anonymous):

duhh . butt how youu qet tht ! ? i need workk ppl

OpenStudy (anonymous):

u gave us two 'equations' ie: 7/x = say 'y' (this 'y' being the number that both of these equal) and also x/7= y ( 'y' being the number that is produced - we can call both these y because u told us they are the same) now we have 2 simultaneous equations: ie: 7/x=y x/7=y there is only one number x which satisfies both of these eqations ie: +7 and -7 as anyhitng else would increas one and decreas the other ... and zero would return a not so good result (u cant divide by zero) also sorry there are two solns for this ... +7 and -7